the population of  a country increased in 10 years at an average rate of 5% per annum. if the  average rate of increase during the first 6 years is 3% p.a. find the average rate of increase in the next 4 years

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

let initial population is P

if for the last 4 years rate is r% then

 P(1+5/100)10= P(1+3/100)6 x (1+r/100)4

take log

10ln1.05 = 6ln1.03 + 4ln(1+r/100)

solving ln(1+r/100)=0.0776

(1+r/100)=1.08

r=8 answer

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..